### v1

# Generating Temporal Networks with the Ascona Model

Abstract
We introduce a new sampling method for continuous-time temporal networks based on queueing processes. In particular, we focus on a Markovian version of the model where the links between nodes are Poisson distributed in time and have exponential duration. We highlight the stochastic properties of these temporal structures and leverage them to design synthetic temporal networks with a controllable level of smoothness, which follow patterns relevant for the validation and interpretation of methods for community, scale, change-point, and periodicity detection. Additionally, we show that imposing assortativity constraints on the samples leads to a continuous-time generalization of stochastic block models. Finally, we describe how variations of the model can be used to sample alternative types of structure and temporal networks, especially discrete-time ones.
1 Introduction
Temporal (or dynamic) networks/graphs are extensions of graphs that incorporate temporal information. Various mathematical objects can be considered to be temporal networks, and different formalisms have been proposed to describe them (e.g., [casteigts_time-varying_2012, holme_temporal_2012, latapy_stream_2018]). These objects can be classified on a spectrum of increasing complexity based on the granularity of the encoded temporal information [rossetti_social_2015, rossetti_community_2018]. On one side, we have static graphs, which aggregate in one structure all temporal information; on the other side, there are stream graphs and link streams that avoid any aggregation. Somewhere in between the two sides lie snapshot networks and discrete-time stream graphs/link streams.
Temporal networks are relevant in modeling many complex systems [masuda_guide_2016], and, as for their static counterparts, understanding their communities/clusters is a fundamental problem in network science. The community structure evolution is typically characterized by archetypal structural changes [cazabet_challenges_2023], called events, or combinations thereof [failla_describing_2024]. Therefore, sampling methods that allow to model these archetypes are of paramount importance for the generation of quality synthetic data with controlled ground-truth. One needs these synthetic networks, for instance, for the validation and interpretation of algorithms tasked for community, scale, change-point, and periodicity detection. A review of the main existing methods, particularly relevant for community detection benchmarks, can be found in [cazabet_evaluating_2021].
For some applications, one would also want to sample networks that are less artificial and instead are closer to real-world datasets. Recently proposed techniques in this direction (e.g., [gauvin_randomized_2022, longa_efficient_2022, presigny_building_2021]) aim at producing surrogate temporal graphs, structures that maintain some or most properties of a reference (often real-world) temporal network. However, these methods do not provide sufficient control to design a controlled experimental setup, as they bypass the explicit formulation of a modeling mechanism by relying on a pre-existing structure.
Agent-based models (e.g., [perra_activity_2012, snijders_introduction_2010]) are an alternative option. They leverage control of microscopic dynamics of a complex system and combine them so that, in the aggregate, complex phenomena emerge. However, often, the understanding of these models relies heavily on simulations, as the underlying dynamics are complex, leading to mathematical intractability, hence limited stochastic and statistical control.
Most of the temporal network literature, especially on sampling methods, has focused on snapshot networks or discrete-time structures, leaving out the more complex continuous-time ones (a notable exception being [bovet_flow_2022]). The main contribution of this article is the introduction of a sampling method for continuous-time temporal networks, and the explanation of how to use it to sample data reflecting the above-mentioned event patterns, plus a continuous-time generalization of the stochastic block model (SBM) [abbe_community_2018].
The model works in two fundamental steps: in the first, we sample when nodes are connected with a link, and in the second, we sample which nodes to assign to each link sampled in the previous step. Separating the temporal from the connectivity aspect offers benefits in terms of implementation and clarity, while not sacrificing expressivity. This separation mechanism also helps to structure the mathematical analysis of a temporal network (model), as the two aspects can be considered orthogonal to each other, hence treated separately, as in [barucca_maximum_2025]. Connectivity can exhibit almost arbitrary levels of complexity and is responsible for various structural aspects such as assortativity, disassortativity, and nestedness. A complete analysis of its properties from a stochastic and statistical perspective is therefore challenging. The temporal aspect appears to be simpler; however, surprisingly, its probabilistic and statistical analysis has been neglected. This work addresses this gap: its key idea is to describe the appearance and disappearance of nodes’ connections in terms of queueing processes, and then focus on the evolution of the total number of active connections as a function of time to derive probabilistic and statistical results.
Queueing processes were originally introduced to describe the flow of customers (jobs, particles, calls) arriving at a service facility. It is possible to approximate the evolution of most queueing systems with Markov chains. The advantage of these is that it is possible to calculate the stationary distribution of a chain by solving a linear system of equations called the balance conditions [walrand_introduction_1988].
The structure of this work is as follows. In Section˜2 we introduce the notation and describe the objects studied in the article. It is divided into two parts: the first is devoted to formalisms of temporal networks, the second to queueing processes. In Section˜3 we introduce the Ascona model, a temporal networks model based on queueing theory, and then use a Markovian parametrization of it to generate link streams with desired temporal evolutions. Based on this Markovian framework, we describe a link stream generalization of SBMs, and compare it to their currently most popular temporal extensions, dynamic SBMs. We conclude the section with a description of variations of the Ascona model. The end of the article, Section˜4, contains a discussion of the presented results and highlights possible future research directions stemming from this work.
2 Theoretical Framework
In this section, we describe the notation and objects we need to introduce the Ascona model. In the first part, we describe various formalisms of temporal networks following the framework presented in [latapy_stream_2018], and in the second part, general queueing processes, focusing on the $M/M/\infty$ queue, also known as the ample-server model [gross_fundamentals_2008, Ch. 2.7].
2.1 Temporal Networks
We model a dynamic network within a time domain $T$ (either an interval of $\mathbb{R}$ for continuous time or a finite subset of $\mathbb{Z}$ for discrete time) and over a finite node set $V$. In the most general setup, stream graphs, $V$ is not fixed in time. However, for many, if not most applications, a disappearing node is equivalent to a node with zero degree that never disappears. Therefore, in Section˜3, we focus on the goal of producing a sampling method for link streams, where $V$ is constant. In Section˜4, we are going to briefly discuss how one could extend the model to generate stream graphs.
2.1.1 Link Streams
The link stream is a temporal network formalism that concentrates on the timing of interactions. There are two common variants.
General Model
A link stream is a triplet
$L=(T,V,E),\qquad E\subseteq\bigl\{[\alpha,\omega)\times\{\{u,v\}\}:\alpha\leqslant\omega,u\neq v\bigr\},$ (2.1)
so that $([\alpha,\omega),\{u,v\})\in E$ means the link $\{u,v\}$ is present for all $t\in[\alpha,\omega)$.
For each $t\in T$ the instantaneous graph is
$G_{t}=(V,E_{t}),\quad E_{t}\coloneqq\bigl\{\{u,v\}:(t,\{u,v\})\in E\bigr\}.$ (2.2)
Given a window $I=[a,b]\subseteq T$, the unweighted footprint is the static graph
$G_{I}\coloneqq\left(V,\bigl\{\{u,v\}:\exists t\in I\text{ with }(t,\{u,v\})\in E\bigr\}\right).$ (2.3)
If $T$ is not discrete, the weighted footprint assigns to each edge the contact time
$w_{I}(u,v)\coloneqq\lambda\left(\{t\in I:(t,\{u,v\})\in E\}\right),$ (2.4)
where $\lambda$ is the Lebesgue measure. If $T$ is discrete, we use the contact count instead.
We say that a stream $S^{\prime}=(T^{\prime},V^{\prime},E^{\prime})$ is a sub-stream of $S=(T,V,E)$ if, intuitively, $T^{\prime}\subset T,V^{\prime}\subset V,E^{\prime}\subset E$.
Instantaneous Contacts
The second variant is a special case of a link stream where links have 0-duration. A link stream with instantaneous events is a triple
$L=(T,V,E),\qquad E\subseteq T\times\binom{V}{2}$, (2.5)
where $(t,\{u,v\})\in E$ records a contact of $\{u,v\}$ at time $t$.
2.1.2 Snapshot Networks
The most common framework to incorporate temporal information in networks consists of splitting the time domain into slices and then building a graph, called a snapshot, for each time slice: its nodes and links represent the interactions that occurred during this time slice. Consequently, snapshot networks consist of a discrete set of graphs and are structurally similar to the above-mentioned discrete-time streams.
The length of each slice is determined based on a compromise: one needs slices large enough to ensure that each snapshot captures significant structural information, but large slices lead to losses of temporal information, since all interactions within the same slice are merged. In addition, several or even varying slice durations may be relevant [latapy_stream_2018].
The above-described models are all natural generalizations of simple graphs to the temporal context and will be the objects of study of this article. One may enrich them by attaching attributes to links or nodes, without changing the underlying $(T,V,E)$ structure, and obtain natural extensions of multiedges, weighted, and signed networks.
2.2 The $M/M/\infty$ Queue
Queues were originally introduced to model the flow of customers arriving at a service facility. They are modeled stochastically as birth-death Markov chains keeping track of the number of elements in the system. The basic ingredients of a queue model are:
  • an arrival process $A$ (typically described by interarrival times),
  • a service-time distribution $S$,
  • a number $c\in\mathbb{N}\cup\{\infty\}$ of parallel servers,
  • a system capacity $K$ that limits the number of customers that can be present,
  • $\bar{K}$ which is the maximal number of elements entering the system,
  • a service discipline $D$ that specifies how customers are selected for service (e.g., first-come, first-served)
Kendall’s notation, originating from [kendall_stochastic_1953], concisely encodes these features as
$A/S/c/K/\bar{K}/D.$ (2.6)
Common defaults are $\infty, K¯=\bar{K}=\infty$, and $D=\text{first-come, first-served (FCFS)}$, in which case one writes simply $A/S/c$.
We are mainly interested in one specific queue, the $M/M/\infty$ one. The ‘$M$’ is a standard symbol in Kendall notation indicating Markovian processes; therefore, the model can be interpreted as a birth (or immigration)-death process where arrivals form a Poisson process with rate $\lambda>0$, and deaths are i.i.d. exponential with rate $\mu>0$ [asmussen_applied_2003]. There are countably infinitely many servers, hence no waiting occurs: each arrival immediately starts the death process.
Let $M(t)$ denote the number of customers in the system at time $t\geq 0$. Then $\{M(t)\}_{t\geq 0}$ is a continuous–time birth–death process on $\mathbb{N}$ with birth rates
$q_{m,m+1}=\lambda\quad\text{for all }m\geq 0,$ (2.7)
and death rates
$q_{m,m-1}=m\mu\quad\text{for }m\geq 1,$ (2.8)
since each of the $m$ customers present departs at rate $\mu$ independently.
The transition rate matrix of the process is
$Q=\begin{pmatrix}-\lambda&\lambda\\ \mu&-(\mu+\lambda)&\lambda\\ &2\mu&-(2\mu+\lambda)&\lambda\\ &&3\mu&-(3\mu+\lambda)&\ddots\\ &&&\ddots&\ddots\end{pmatrix}$. (2.9)
Writing $p_{m}(t)=\mathbb{P}(M(t)=m)$, the corresponding Kolmogorov forward equations are
$p_{0}^{\prime}(t)$ $=-\lambda p_{0}(t)+\mu p_{1}(t),$ (2.10)
$p_{m}^{\prime}(t)$ $=\lambda p_{m-1}(t)-(\lambda+n\mu)p_{m}(t)+(n+1)\mu p_{m+1}(t),\qquad m\geqslant 1.$ (2.11)
2.2.1 Process Evolution
It is known (e.g., [robert_stochastic_2013, Ch. 6]) that, assuming that $M(0)=0$, $M(t)$ is Poisson with a time–dependent mean
$m(t)=\frac{\lambda}{\mu}\left(1-e^{-\mu t}\right);$ (2.12)
hence, for every $j\in\mathbb{N}$, the transition probabilities from 0 are
$\mathbb{P}\{M(t)=j\mid M(0)=0\}=\exp\!\left(-\frac{\lambda}{\mu}\left(1-e^{-\mu t}\right)\right)\frac{\left(\frac{\lambda}{\mu}(1-e^{-\mu t})\right)^{j}}{j!}.$ (2.13)
We can derive this fact by observing that the probability generating function $G(z,t)\coloneqq\sum_{m\geq 0}p_{m}(t)z^{m}$ solves
$\frac{\partial G(z,t)}{\partial t}=(z-1)\lambda G(z,t)-(z-1)\mu\frac{\partial G(z,t)}{\partial z}$ (2.14)
with $G(z,0)\equiv 1$; yielding
$G(z,t)=\exp\!\Big(m(t)(z-1)\Big),$ (2.15)
the probability generating function of a $\mathrm{Poisson}(m(t))$ law.
This gives
$\mathbb{E}[M(t)]=\frac{\lambda}{\mu}\left(1-e^{-\mu t}\right),$ (2.16)
and
$\mathrm{Var}\!\left(M(t)\right)=\frac{\lambda}{\mu}\left(1-e^{-\mu t}\right).$ (2.17)
If we stop sampling elements after a certain time $t_{e}$, then the process is truncated, and the mean follows the equation
$m(t)=\begin{cases}0,&t<t_{s},\\[4.0pt] \lambda\mu\left(1-e^{-(t-t_{s})/\mu}\right),&t_{s}\leq t\leq t_{e},\\[4.0pt] \lambda\mu\left(e^{-(t-t_{e})/\mu}-e^{-(t-t_{s})/\mu}\right),&t>t_{e}.\end{cases}$ (2.18)
We display these properties in Figure˜1.
Figure 1: Comparison of analytical formulas and trajectories of 100 $M/M/\infty$ queue samples. In the top plot, we compare the average density of the sample as a function of time with the analytical formula Eq. (2.18). The fluctuations of the trajectories are in the square root order of the mean profile, as predicted analytically. The comparison of this analytical formula for the standard deviation with the sample one is displayed in the bottom plot.
2.2.2 Stationary Regime
The $M/M/\infty$ chain is always irreducible and positive recurrent; this is in contrast with finite servers queues $M/M/c$, where $\frac{\lambda}{c\mu}<1$ is required for ergodicity. So, if we do not stop sampling after a certain time-point $t_{e}$, we converge towards a stationary behavior.
One can get the asymptotic behavior by noticing that $m(t)\to\lambda/\mu$ and that the Poisson family is closed under weak limits, hence
$M(t)\ \xrightarrow[]{d}\ \mathrm{Poisson}\!\left(\frac{\lambda}{\mu}\right).$ (2.19)
Alternatively, the stationary probabilities $\{\pi_{m}\}_{m\geq 0}$ satisfy the detailed balance equations
$\pi_{m}\lambda=\pi_{m+1}(m+1)\mu\quad\text{for }m\geq 0,$ (2.20)
expressing equilibrium between the birth and death processes of the queue. Solving recursively gives
$\pi_{m}=\pi_{0}\frac{(\lambda/\mu)^{m}}{m!}.$ (2.21)
Normalizing with $\sum_{m=0}^{\infty}\pi_{m}=1$ yields
$\pi_{m}=\mathbb{P}\{M(\infty)=m\}=e^{-\rho}\frac{\rho^{m}}{m!},\qquad m=0,1,2,\dots$ (2.22)
i.e., the stationary distribution is $\mathrm{Poisson}(\rho)$ with mean $\rho\coloneqq\frac{\lambda}{\mu}$.
Consequently,
$\mathbb{E}[M(\infty)]=\operatorname{Var}(M(\infty))=\frac{\lambda}{\mu}.$ (2.23)
One could also recover $\mathbb{E}[M(\infty)]=\frac{\lambda}{\mu}$ directly from Little’s law (see [little_proof_1961] and [walrand_introduction_1988, Thm. 3.4.1]).
We now derive an estimate for the time $t^{\star}$ at which we can assume that $m(t)$ is close enough to the stationary value. We use this estimate later in Section˜3.4 and Section˜3.5.
2.2.3 Head of the Queue
We consider an $M/M/\infty$ queue that starts empty at time $t=0$, with arrival rate $\lambda>0$ (Poisson arrivals) and i.i.d. service times $S\sim\mathrm{Exp}(\mu)$ of mean $\mu>0$.
One can heuristically think of being at steady state as if the system had “forgotten” its empty start, so the mean occupancy is close to
$m_{\infty}\coloneqq\frac{\lambda}{\mu}.$ (2.24)
We define the (deterministic) distance from stationarity at time $t$ as
$\delta(t)\coloneqq m_{\infty}-m(t)=\frac{\lambda}{\mu}e^{-\mu t}.$ (2.25)
Because $M(t)$ is Poisson with mean $m(t)$, its standard deviation at time $t$ is
$\sigma(t)=\sqrt{\mathrm{Var}[M(t)]}=\sqrt{m(t)}=\sqrt{\frac{\lambda}{\mu}\left(1-e^{-\mu t}\right)}.$ (2.26)
We now formalize the idea that the system has become “indistinguishable from stationarity” as the first time $t^{\star}$ when the initial bias $\delta(t)$ is no larger than the typical intrinsic fluctuation scale $\sigma(t)$. Concretely, we are interested in the crossover time $t^{\star}$ defined by
$\delta(t^{\star})=\sigma(t^{\star}).$ (2.27)
The computations in Appendix˜A show that
$t^{\star}=-\frac{1}{\mu}\ln\!\left(\frac{-1+\sqrt{1+4\frac{\lambda}{\mu}}}{2\frac{\lambda}{\mu}}\right),$ (2.28)
and that if $\frac{\lambda}{\mu}$ is moderately large, we can approximate it as
$t^{\star}\approx\frac{1}{2\mu}\ln(\frac{\lambda}{\mu}).$ (2.29)
In the case of $N$ queue samples, the standard deviation of the sample mean is $\sqrt{\frac{m(t)}{N}}$, so it takes more time to establish closeness to stationarity. The computation is equivalent to that of the single sample case. This heuristic argument can be made rigorous in terms of the total variation distance, but a more refined estimation based on mixing times [levin_markov_2017] is possible, allowing one to control closeness to stationarity to an arbitrary level of precision.
2.2.4 Number of Elements in the Queue
The number of elements that join the queue in a time interval $t_{1},t_{2}]$ is, by design, Poisson distributed with average $\lambda(t_{2}-t_{1})$. The number of elements in the queue in the same interval is slightly different. In addition to new elements joining in that interval, we need to count how many started before and ended after $t_{1}$. This amounts approximately to the sum of two independent random variables, one Poisson with average $\lambda(t_{2}-t_{1})$ and the other Poisson with average $m(t_{1})$, which by superposition is again Poisson with average $\lambda(t_{2}-t_{1})+m(t_{1})$. In the stationary regime, this becomes a Poisson random variable with average $\lambda(t_{2}-t_{1})+\lambda/\mu$. We use these facts later in Section˜3.5.
3 The Ascona Model and EDEDE Structures
We design a link streams model based on queueing processes. In the absence of a previously established name, we will refer to it as the Ascona model/method for clarity of the presentation and convenience.
3.1 The Ascona Model
The method relies on the sampling of fundamental sub-stream units called queue blocks sharing the same vertex set $V$ and time domain $T$. Each block links set $E$ is determined in two separate steps: in the first, we create a timeline of the links belonging to it, and in the second, we assign each link to a pair of nodes. The global network is then obtained by combining queue blocks sequentially and in parallel. We first start by describing the creation of a single queue block.
3.1.1 Temporal Structure
The number of links in a queue block, their start and ending times are sampled based on a queueing process $A/S/c/K/\bar{K}/D$ in the interval $0,t_{e}]$, plus the tail obtained by stopping the admission of new elements in the queue after $t_{e}$. The choice of the queue parameters can be based on ad hoc properties of the repartition in time of links, or prior information about it observed in a reference network. The same goes for the choice of which section of the queue to retain (e.g., keep only the head, the tail, or the stationary part of the queue as discussed later in Section˜3.4) and where to situate the block in the time domain (i.e., applying a translation factor to all elements of the queue). In this work, we deal exclusively with ad hoc selection and placements of queue blocks; however, one could imagine randomizing this process in various ways.
The maximal possible number of links at each time is $\binom{N}{2}$, however we can still work with queues with $K>\binom{N}{2}$ if we have a mechanism to stop sampling links when the queue is congested, as we explain later. For instance, in Section˜3.2, we discuss a fundamental queue parametrization, which assumes Markovian arrival and service times, plus infinitely many servers.
3.1.2 Connectivity Structure
After sampling the number of links and their starting and ending times, we need to assign each link to a pair of nodes. We do it stochastically by sampling according to a probability distribution on $V\times V$. We can encode such a distribution in an $N\times N$ matrix whose entries add up to one. We will refer to it as the connectivity probability matrix.
Alternatively, it is often convenient to split the selection process into two. In the first step, a node is picked up randomly according to a probability distribution whose support is a subset of $V$. For instance, a convenient choice would be the uniform distribution, which has maximal entropy and therefore can be interpreted as uninformative from a Bayesian perspective. One could bias this choice based on prior knowledge of (potentially uneven) nodes’ activity, a crucial point in [perra_activity_2012]. In the second step, we sample the second node based on a probability distribution on $V$ dependent on the first node. We can encode these distributions in an $N\times N$ row-stochastic matrix. We will refer to it as the connectivity conditional probability matrix, as the $(i,j)$-th entry corresponds to the probability of selecting $v_{j}$ for a link, given that the first sampled node was $v_{i}$.
In both cases, we impose that the diagonal elements of the connectivity matrices are all equal to zero to avoid self-links.
We show a concrete example of two dynamic networks sampled with the Ascona method in Figure˜2. The two samples share the same temporal structure but have different connectivity. The shared temporal structure is sampled based on an $M/M/\infty$ queue, whose construction is represented in Figure˜3. We discuss in more detail this fundamental parametrization of the Ascona model in Section˜3.2.
3.1.3 Conflicts
A conflict arises in the sampling mechanism whenever we want to resample an already present link. A possible cause for such type of conflicts is that the queue is congested and close to saturation. We can estimate if a model is prone to this problem by verifying its likelihood of getting close to the saturation value $\binom{N}{2}$. If a queue has a stationary value much lower than that value, and the variance of its trajectories is limited, this likelihood will be small, and conflicts due to congestion will occur rarely.
Another cause for conflicts is connectivity constraints. For instance, consider the situation of a node $v_{i}$ already connected to a node $v_{j}$, and a new link originates from $v_{i}$. If, according to the probabilistic connectivity matrix, $v_{i}$ connects exclusively with a node $v_{j}$, the new link will generate a conflict. The number of issues hence increases if the connectivity constraints increase the probability of resampling the same pairs. We can gauge how problematic a certain connectivity probability matrix is by computing its Shannon entropy. Similarly, to gauge a connectivity conditional probability matrix we can compute its conditional entropy conditioned on the sampling vector; analogously to [koovely_evolution_2025], where the probabilistic matrix is a transition operator between nodes instead of a connectivity one.
If a conflict arises, it means that after a link $(v_{1},v_{2},s_{1},e_{1})$, we sampled a link $(v_{1},v_{2},s_{2},e_{2})$ with $s_{2}<e_{2}$. Three possible solutions are:
  • merge the two links in one $(v_{1},v_{2},s_{1},\max(e_{1},e_{2}))$,
  • discard all elements of the second link,
  • resample the pair of nodes but keep the interval $(s_{2},e_{2})$ as basis for the link.
All suggested solutions lead to small discrepancies from the desired temporal or connectivity structure. One might choose a solution to mitigate discrepancies from one or the other aspect, or based on practical considerations: for instance, if a queue is saturated, resampling nodes won’t work.
Notice that the likelihood of errors decreases as the number of nodes increases. The presence of an error does not increase the probability of another error occurring, so they do not propagate, and actually, there is a self-adjustment mechanism: if the appearance of a link causes a conflict, then the discrepancy caused by it ends before the problematic link’s disappearance.
Figure 2: Example of queue blocks $L^{\prime}$ and $L^{\prime\prime}$ sharing the same links’ repartition in time (sampled from a $M/M/\infty$ process), but different connectivity structures. For each of the two sampled link streams, we display the adjacency matrix of the instantaneous graphs $G_{t}^{\prime}$ and $G_{t}^{\prime\prime}$ at three different times $(t\in\{20,50,80\})$. On top, we show the ones of $L^{\prime}$, that has a constant two-block connectivity structure. On the bottom side, we show the ones of $L^{\prime\prime}$, that does not have any particular structure: all pairs of nodes have the same probability to be connected, hence this could be considered a link stream version of an ER graph.
3.2 EDEDE Structures
An EDEDE structure is a link stream generated based on the following inputs: a starting time $t_{s}$, an ending time $t_{e}$, an activity rate $\lambda$, and an interaction rate $\mu$. First, we impose that the starting times of each event are distanced by i.i.d. exponential random variables with parameter $\lambda$, or equivalently that the starting times are Poisson distributed with average $\lambda(t_{e}-t_{s})$. Additionally, we require that all events have i.i.d. exponential duration with parameter $\mu$.
We therefore have Exponential-Duration Events Distanced Exponentially, hence the acronym EDEDE.
We can conveniently sample such structures with the Ascona method. The underlying queue model is $M/M/\binom{N}{2}/\binom{N}{2}/\infty/\text{FCFS}$ [gross_fundamentals_2008, Ch. 2.8], a truncated version of the $M/M/\infty$ queue [robert_stochastic_2013, Ch. 6.7]. However, we can assume to be working in the regime $\binom{N}{2}\gg\lambda\mu$ so that we can consider the distortions due the finite-source constraint as boundary effects and the model is approximately a $M/M/\infty$ model.
Figure 3: Representation of the generation of an EDEDE structure with the Ascona model. (A) The first step consists of drawing i.i.d. exponentially distributed time intervals with rate $\lambda$, or conceptually equivalently, time points based on a Poisson process with rate $\lambda$, over a desired time interval. Each sampled time point corresponds to a starting time $s_{i}$ of a link of the dynamic network; in this example, there are five of them: $s_{1},\dots,s_{5}$. (B) The second step consists of sampling a time interval for each drawn starting point based on i.i.d. exponentially distributed time intervals with rate $\mu$. Each drawn interval describes the duration of the corresponding link. In this example, we end up with five link intervals: $(s_{1},e_{1}),\cdots,(s_{5},e_{5})$. (C) Finally, each link is associated with a pair of vertices. For each link, the first element of the pairs is drawn uniformly at random from the node set. Its interacting partner is drawn randomly based on a distribution encoded in the connectivity conditional probability matrix displayed on the left. In this case, the connectivity matrix encodes a two-block structure with high probability of generating links within each block, and low probability of link generation between blocks. We display node names on a vertical axis on the left of the figure. Each link is represented by a coloured rectangle showing the involved nodes and the time interval when it happens.
From an information-theoretical or Bayesian perspective, an EDEDE structure can be seen as a reference null model that does not assume any knowledge of any correlation between arrival and service times: the exponential distributions are uninformative priors because they have maximum differential entropy among all continuous distributions supported in $0,\infty]$ with the same mean [dowson_maximum-entropy_1973].
However, one might be interested in sampling links based on different distributions, other than exponential ones [barucca_maximum_2025]. In that case, we obtain queues alternative to the $M/M/\infty$ one, whose implementation leads to small modifications of the Ascona sampling method, but lead to structures different from the EDEDE ones. Theoretical properties of these derived models depend on the theoretical properties of these alternative queueing processes; however, some properties like the stationary value are often independent from the Markovian assumptions [bunday_gmr_1980, gross_fundamentals_2008]. See [gross_fundamentals_2008] for a discussion of the consequences of dropping the purely Markovian assumption for the arrival times, in particular allowing restrictions on the support, inhomogeneity of the rate parameter, and batch arrivals.
3.2.1 Degenerate Cases
We can generate link streams with 0-duration links with the Ascona model by setting the service times to be Dirac $\delta$ distributed. In the EDEDE structure case, this reduces to a system with links spawning as a Poisson process.
Instances of discrete-time link stream can be sampled by picking arrival times, and edge durations according to $\mathbb{N}$-supported distributions.
3.3 Smooth Stochastic Blocks
In this part, we describe how to use the Ascona model to sample link streams with evolving community structures. The idea is to design a single queue block where the connectivity follows the desired clusters’ evolution. In this article, we show EDEDE samples exclusively, but the desired effect is achievable almost independently of the underlying queue process by imposing an appropriate connectivity (conditional) probability matrix that evolves as a function of time.
In the example displayed in Figure˜2, we used constant two-block connectivity conditional probability matrices, leading to a constant community structure. In the example displayed in Figure˜5, we first have a four-block structure followed by a two-block structure. The change represents a discontinuity in the progression of the underlying connectivity conditional probability matrix, however preexisting links take a while to disappear after the change, while links associated with the new structure appear gradually, leading to smooth transitions, hence the name “smooth stochastic blocks".
In Section˜3.5, we explain how to aggregate smooth stochastic blocks in order to get snapshots, and after that, we compare them to Dynamic SBMs, the main generalization of SBMs to temporal networks. Now, we showcase the expressivity of the newly introduced smooth SBMs.
3.4 Generating Archetpyes
A fundamental aspect of temporal networks is the understanding of its clusters and how they evolve in time. Similarly to the methods used in [asgari_mosaic_2023] and [cazabet_evaluating_2021], the procedure to generate link stream scenarios follows two steps: first, the experimenter needs to describe the scenario in terms of a queue blocks’ footprint, and secondly, generate the structure with the Ascona model to obtain a temporal network that qualitatively reflects the desired scenario.
In this part, we display the expressivity of the Ascona model by describing how to generate the typical “events" characterizing temporal networks’ evolution listed in [cazabet_challenges_2023]. For elementary versions of such scenarios, a single EDEDE queue block is sufficient. In order to design some of these events, one can focus exclusively on the temporal aspect of the model.
Birth/Death
In a birth event, a new community appears. Conversely, in a death event, an existing community vanishes. One can model the birth of a community by either taking the head or the time-reversed tail of an EDEDE queue block. In the first case, the growth is not necessarily monotonic and has roughly a concave shape. In the second case, the growth is monotonic and has approximately a convex shape. Conversely, one can model the death of a community by taking the time-reverse of a birth event.
Continuity
One can model a stable community by taking a queue block in its stationary regime.
Change of intensity
One event, not listed among the ones in [cazabet_challenges_2023], is the one where a stable community experiences an increase/decrease in the number of interactions. We can model this situation with an EDEDE queue block by imposing that the parameters $\lambda$ and $\mu$ of the underlying $M/M/\infty$ process change at a switching time point $t_{sw}$. The underlying process remains Poissonian, and its average is
$m(t)=\begin{cases}0,&t<t_{s},\\[6.0pt] \frac{\lambda_{1}}{\mu_{1}}\!\left(1-e^{-\mu_{1}(t-t_{s})}\right),&t_{s}\leq t\leq t_{sw},\\[8.0pt] \frac{\lambda_{1}}{\mu_{1}}\!\left(e^{-\mu_{1}(t-t_{sw})}-e^{-\mu_{1}(t-t_{s})}\right)\;+\;\frac{\lambda_{2}}{\mu_{2}}\!\left(1-e^{-\mu_{2}(t-t_{sw})}\right),&t_{sw}\leq t\leq t_{e},\\[8.0pt] \frac{\lambda_{1}}{\mu_{1}}\!\left(e^{-\mu_{1}(t-t_{sw})}-e^{-\mu_{1}(t-t_{s})}\right)\;+\;\frac{\lambda_{2}}{\mu_{2}}\!\left(e^{-\mu_{2}(t-t_{e})}-e^{-\mu_{2}(t-t_{sw})}\right),&t>t_{e}\end{cases}$
We display in Figure˜4 the quality of this formula with simulations of $M/M/\infty$ queue experiencing an increase in activity due to changing parameters.
Figure 4: Trajectories of 100 samples of a $M/M/\infty$ queue with changing parameters at time $t_{sw}=100s$. Before $t_{sw}$, the rate of arrivals is $\lambda_{1}=2.5$ and the service time rate $\mu_{1}=0.1$; after $t_{sw}$, they are equal to $\lambda_{2}=5$, and $\mu_{2}=20$. We compare the sample mean with the analytical mean from Eq. (3.1) until the time $t_{e}=200$ when we stop new elements from joining the queue.
More complex temporal scenarios require a chain of queue blocks.
Resurgence
Resurgency can be thought of as a combined event where a queue block vanishes for a period, then reappears without particular perturbations.
Other events, the ones that have more to do with how communities interact, can be designed by intervention on the connectivity aspects of a queue block.
Merge/Split
A merge event, as the name says, indicates that multiple existing clusters of nodes merge into one. Conversely, in a split, an existing cluster divides into multiple smaller clusters. One can easily engineer such events with a queue block that follows a certain connectivity matrix before the merge/split event, and a new one after.
Growth/Contraction
The growth and contraction events are similar to the previous case. In a growth event, we observe an existing community acquire new nodes; in a contraction, a community loses some nodes, but does not disappear completely. In these cases, instead of two interacting communities, one community interacts with a set of unstructured, lowly connected nodes.
One can notice that this procedure leads to controllably smooth transitions, balancing stability in time of the links with consistency with the connectivity structures, a known problematic aspect of sampling temporal networks [cazabet_evaluating_2021, Section 3.2.1].
We display in Figure˜5 the generation of many of these events with a single EDEDE block, encoding the evolution of a link stream from a four-block to a two-block structure. We show as well that by slicing the time domain and aggregating the information over the slices, we get a snapshot equivalent of this link stream. We discuss this procedure in more detail in the next part.
Figure 5: Ascona model sample displaying various events. (a) Birth event of four communities. (b) In the stationary regime, there is a stable configuration with four fully developed communities. (c) At time 100, there is a merge event leading to the formation of two bigger communities. (d) After time 200, the queue decays, and the two communities vanish. (A) Weighted footprint of the network of the interval $[40s,60s]$ displaying clearly the four-block structure before the merge. (B) Weighted footprint of the interval $[90s,110s]$ displaying the transition between the four and the two-block structure happening at time $100s$. (C) Weighted footprint of the interval $[140s,160s]$ displaying the two-block structure after the merge. The colour of each footprint entry depends on the weight from Eq. (2.4). Intuitively, warmer colours indicate higher weights.
3.5 Generating Snapshot Networks
As mentioned previously, the majority of datasets and methods in temporal networks theory has been designed primarily (and often exclusively) with snapshot networks in mind. We therefore describe a procedure for generating snapshot networks from an EDEDE sample in a controlled way. The strategy is to discard the head of the queue based on the estimated crossover times (Eq. (2.28) or its approximation Eq. (2.29)) presented in Section˜2.2.3, and then slice the time domain and aggregate the links in unweighted footprints (see Eq. (2.3)) which leads to “uniform" snapshots.
We suggest two uniformity criteria for generating homogeneous snapshots. The first criterion is to try to produce snapshots from time intervals with the same length $\Delta t$, so that they contain the same average number of spawning links. Let this number be $K$; then, from Section˜2.2.4, we know that in the stationary regime,
$\lambda\Delta t\overset{!}{=}K,$ (3.2)
which leads to
$\Delta t=\frac{K}{\lambda}.$ (3.3)
Another uniformity criterion is to impose $K$ to be the total number of active links, on average, in each snapshot. In this case, always from Section˜2.2.4, we know that, in the stationary regime, we have
$\lambda\Delta t+\frac{\lambda}{\mu}\overset{!}{=}K,$ (3.4)
which leads to
$\Delta t=\frac{K}{\lambda}-\frac{1}{\mu}.$ (3.5)
In both cases, one could increase the overlap between subsequent slices to achieve a higher level of smoothness. Notice that if one wants to exert control on the number of distinct links in every slice, then a correction factor for Eq. (3.3) and Eq. (3.5) is needed to take into account the probability of links appearing multiple times in a certain $\Delta t$.
3.5.1 Comparison with Dynamic SBM
If we apply the above-presented procedure to smooth SBM link streams, we obtain snapshots with block structures. This is reminiscent of the Dynamic SBM model, which has emerged as a popular extension of its static network counterpart and is used, for instance, for testing change-point detection methods (e.g., in [bhattacharjee_change_2020, sulem_graph_2024]). In dynamic SBMs, each snapshot is sampled according to a static SBM distribution whose parametrization can change in time. In terms of expected number of edges and its variance, smooth SBMs and dynamic SBMs are similar because of the similarity between Poisson and Binomial distributions. The two methods differ in terms of smoothness: dynamic SBM snapshots have no mechanism to enforce smoothness, while the smoothness of the underlying smooth SBM link stream is inherited by the derived SBM snapshots.
4 Conclusions
This article focused on sampling link streams with non-zero duration links, but throughout the work, we highlighted steps where minor additions or modifications would lead to different possibilities, such as multi/self/weighted/signed links. Additionally, we discussed how to obtain other types of temporal graphs: discrete-time link streams, 0-duration link ones, and snapshot networks. Graph streams are more complex than link streams, and one requires an extra ingredient to sample them compared to link streams: a sampling mechanism for the temporal presence of each node.
In Section˜3.3, we introduced smooth SBM, a specific variation of the Ascona model that generates link streams with community structures, and compared it with the dynamic SBM model. Similarly, one could design alternative (conditional) probability connectivity matrices to generate structures with different properties, e.g., nestedness, disassortativity, and structurelessness. In Section˜3.4, we have seen how to generate archetypical temporal structures with the smooth SBMs. As discussed in [failla_describing_2024], hybridization of these simple archetypes leads to more complex structures. It is realistic to imagine generating hybrids by combining several queue blocks, and this is a possible future direction of this work, raising the question of how expressive the Ascona model really is.
Aside from these positive aspects of expressivity and flexibility, a big advantage of this article’s approach is that it translates temporal graphs’ problems to problems of time series and stochastic processes that are well studied and understood, with a rich body of theoretical and practical results ready to be exploited. For instance, [walrand_introduction_1988] discusses various aspects of statistical analysis on queues. $M/M/\infty$ models, the foundation of the EDEDE structures, are particularly nicely behaved from a statistical perspective: for instance, the MLE estimation is equivalent to the method of moments when fitting the exponential distributions of the model. Other queue models would require the use of simulations, but some are remarkably amenable to analyze (e.g., see [eick_physics_1993]).
Future work should address aspects of inferential statistics: given a reference link stream, find the Ascona model parametrization that fits the data in the best way possible. Recent works on maximum entropy approaches to temporal networks [barucca_maximum_2025, clemente_temporal_2024] could pave the way in this direction and go beyond the purely Markovian framework.
Material
A GitHub project containing the code used for the generation of all plots is available [^2] https://github.com/samuelkoovely/EDEDE_project.
Appendix A Head of the Queue: Computations
Here we derive the crossover time $t^{*}$ as defined in Eq. (2.27).
Substituting Eq. (2.25) and Eq. (2.26) into Eq. (2.27) gives
$\frac{\lambda}{\mu}e^{-\mu t^{\star}}=\sqrt{\frac{\lambda}{\mu}\left(1-e^{-\mu t^{\star}}\right)}.$ (A.1)
Let
$\rho\coloneqq\frac{\lambda}{\mu},\qquad x\coloneqq e^{-\mu t^{\star}},\quad\text{so that}\quad t^{\star}=-\frac{\ln x}{\mu}.$ (A.2)
Then Eq. (A.1) becomes
$\rho x=\sqrt{\rho\left(1-x\right)}.$ (A.3)
Squaring both sides yields
$\rho^{2}x^{2}=\rho(1-x)\quad\Longrightarrow\quad\rho x^{2}+x-1=0.$ (A.4)
Equation Eq. (A.4) is quadratic in $x$ with solutions
$x=\frac{-1\pm\sqrt{1+4\rho}}{2\rho}.$ (A.5)
Because $x=e^{-t^{\star}/\mu}$ must be positive, we take the ‘++’ root:
$x=\frac{-1+\sqrt{1+4\rho}}{2\rho}.$ (A.6)
Finally,
$t^{\star}=-\frac{1}{\mu}\ln\!\left(\frac{-1+\sqrt{1+4\frac{\lambda}{\mu}}}{2\frac{\lambda}{\mu}}\right).$ (A.7)
Equation Eq. (A.7) is an exact expression for the crossover time $t^{\star}$ at which the deterministic distance from stationarity falls to the same order as the instantaneous stochastic fluctuations.
We can approximate this value when the steady-state mean load $\rho$ is large enough ($\rho\gg 1$). In that case we can simplify Eq. (A.6) using $1+\sqrt{1+4\rho}\approx 2\sqrt{\rho}$:
$x\approx\frac{2\sqrt{\rho}}{2\rho}=\frac{1}{\sqrt{\rho}},\qquad\text{so}\qquad t^{\star}\approx-\frac{1}{\mu}\ln\left(\frac{1}{\sqrt{\rho}}\right)=\frac{1}{2\mu}\ln(\rho).$ (A.8)
Thus, for moderately large $\frac{\lambda}{\mu}$, an accurate rule of thumb is
$t^{\star}\approx\frac{1}{2\mu}\ln(\frac{\lambda}{\mu}).$ (A.9)
Acknowledgments
Y. Asgari and G. Vaccario suggested articles relevant to the subject, while A. Bovet and Z. Fang’s comments helped me to improve this manuscript. C. Casati helped in finding the catchy palindromic acronym EDEDE. My gratitude goes to all of them.